Let x be 2. The firm has an average collection period of 34 days. Current practice is to factor all receivables immediately at a discount rate of (1+0.1*x)%. Assume that default is extremely unlikely. What is the effective annual interest rate on this arrangement?

A) 11.39 percent

B) 12.61 percent

C) 13.84 percent

D) 15.08 percent

E) 16.34 percent
Answer #1

Answer is 13.84%

Discount Rate = (1 + 0.1*x)%
Discount Rate = (1 + 0.1*2)%
Discount Rate = 1.20%

Average Collection Period = 34 days

Effective Interest Rate = [1 + Discount Rate / (1 - Discount Rate)]^[365 / Average Collection Period] - 1
Effective Interest Rate = [1 + 0.0120 / (1 - 0.0120)]^[365 / 34] - 1
Effective Interest Rate = [1 + 0.0120 / 0.9880]^[365 / 34] - 1
Effective Interest Rate = 1.012146^10.735294 - 1
Effective Interest Rate = 1.1384 - 1
Effective Interest Rate = 0.1384 or 13.84%